# RG approach to the inviscid limit for shell models of turbulence

### v1

@@ 6.1 RG formalism
The RG formalism is defined by the symmetries and couplings of the ideal IBVP. Similarly to the dyadic model, the symmetries are the time and space scalings. The time scaling symmetry has the same form (2.6). For the boundary condition (6.2), the space scaling symmetry is forulated as
$\tilde{u}_{n}(t)=\lambda u_{n+1}(t),\quad\tilde{a}_{n}=\lambda a_{n+1},\quad\tilde{b}(t)=(\lambda b_{0}(t),\,\lambda u_{1}(t)).$ (6.3)
The RG formalism is introduced in the same way as in Section 3; see Definitions 1 and 2. Here Eqs. (3.4) are adapted to the boundary condition (6.2) as
$\tilde{a}_{n}=\lambda a_{n+1},\ n\geq 1;\quad\tilde{b}(t)=\big{(}\lambda b_{0}(t),\,\lambda u_{1}(t)\big{)}.$ (6.4)
The RG operator $\mathcal{R}$ acts in the space of canonical regularizations. Canonical regularizations are flow maps $\Phi$, which are infinitely renormalizable and invariant with respect to the time scaling. Since the RG operator maps $N$-level to $(N+1)$-level canonical regularizations, the vanishing regularization limit is associated with the RG dynamics $\mathcal{R}^{N}[\Phi]$ as $N\to\infty$. Recall that the RG formalism separates the roles of the ideal system and regularization: the RG operator depends on the coupling function (6.1) of the ideal model, while all information on regularization is contained in the flow map $\Phi$.
@@ 6.2 RG attractor
Figures 4(a,b) show evolutions of shell variables $u_{n}(t)$ of the regularized model (3.7) and (6.1) with $(N,J)=(20,3)$ for two initial conditions (4.9) and boundary conditions $b_{-1}(t)=1$ and $b_{0}(t)=2+\sin t$. Figures 4(c,d) show the respective evolutions of the first two variables $(u_{1},u_{2})$ for different cutoff parameters $N=30,\ldots,50$. For the initial condition $\mathrm{IC}_{1}$ (panel c), the figures verify that the regularized solutions converge at pre-blowup times $t\leq t_{b}\approx 3.63$ but diverge at larger times $t>t_{b}$. For the initial condition $\mathrm{IC}_{2}$ (panel d), the regularized solutions diverge at all positive times.
Figure 4: Evolutions of all shell variables $u_{n}(t)$ for the regularized model (3.7) and (6.1) with $(N,J)=(20,3)$ and the initial condition (a) $\mathrm{IC}_{1}$ and (b) $\mathrm{IC}_{2}$. Lower panels show the respective evolutions of the first two variables $(u_{1},u_{2})$, where the cutoff parameter is changed in the range $N=30,\ldots,50$. Here the color changes gradually from red for $N=30$ to blue for $N=50$.
@@
Let us consider regularizations (3.7) with $J=3$ and modified dissipative terms $c_{n}|u_{n}|u_{n}$, which are multiplied by positive coefficients $c_{n}$. We perform $100$ simulations for each $N=40,\ldots,80$ by choosing $c_{n}$ randomly from the interval $[0,3]$. The resulting values $(u_{1},u_{2})$ are presented by black dots in Fig. 5, which form closed curves in the phase space. According to Eq. (3.8), each random choice of regularization corresponds to a specific RG sequence $\mathcal{R}^{N}[\Phi]$ starting from a different initial flow map $\Phi$. Hence, the black dots forming a curve in Fig. 5 determine the RG attractor probed by 100 different initial conditions. Similar curves appear for any pair of shell variables. The one-dimensional asymptotic structure is also apparent in Figs. 4(c,d), where the data from Fig. 5 is added at final times.
@@
Our numerical results suggest that the RG dynamics has a quasiperiodic attractor [22]: it is a closed curve $\Phi^{(\theta)}$ in the space of flow maps parametrized by points of a circle, $\theta\in S^{1}$. This attractor leads to a more sophisticated (than in the fixed-point case) but still universal description of the vanishing regularization limit. Namely, within the basin of attraction, $\Phi\in\mathcal{B}(\Phi^{(\theta)})$, the RG dynamics takes the asymptotic form
@@
where $\Phi^{(\theta_{N})}$ are flow maps from the RG attractor. The phases $\theta_{N}$ depend on the initial flow map $\Phi$, and are related by the circle map $\mathcal{R}_{\theta}:\theta_{N}\mapsto\theta_{N+1}$ induced by the RG dynamics on the attractor. In other words, for any canonical regularization from the basin of attraction, regularized solutions $u(t)=\mathcal{R}^{N}[\Phi]_{t}(a,b)$ approach asymptotically a universal (i.e., independent of regularization) family of solutions $\Phi^{(\theta)}_{t}(a,b)$. The latter is a one-parameter family of solutions of the ideal IBVP.
@@ 6.3 Inviscid limit
Here the parameter $\theta$, which selects a map from the RG attractor, depends on $(N,\beta)$, which, in turn, are functions of viscosity, time, initial and boundary conditions. This dependence implies that the universal relation (6.6) is not valid for finite times, unlike in the case of canonical regularizations (6.5).
Figure 6: (a) Evolutions of the first two variables $(u_{1},u_{2})$ for the viscous model with the second initial condition $\mathrm{IC}_{2}$. Different curves correspond to different viscosities logarithmically spaced in the interval $1.65\times 10^{-10}\leq\nu\leq 10^{-6}$. (b) Terminal points (at time $t=0.5$) of the viscous solutions from the left panel are compared with the data from Fig. 5(b) featuring the quasiperiodic RG attractor.
@@ 7 Chaotic RG dynamics in the Sabra model
The RG formalism for the Sabra model is developed in the same way as in Sections 3 and 6. We learned in Section 3 that the RG operator is defined unambiguously if it is restricted to canonical flow maps, which are maximally symmetric. The latter means that the flow maps are invariant with respect to all symmetries except the space scaling (6.3). In the Sabra model, these are the time scaling and phase symmetries. From now on we assume that all regularizations under consideration are invariant with respect to the phase symmetry (7.2). For a flow map $\Phi$ this condition means that the regularized solutions $u(t)=\Phi_{t}(a,b)$ and $\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b})$ are related by Eq. (7.2). One can see that the viscous model (2.7) is phase invariant, but not invariant to the time scaling.
Figure 7: Absolute values of shell variables $|u_{n}(t)|$ for the regularized Sabra model with $(N,J)=(17,2)$ and the initial conditions (a) $\mathrm{IC}_{1}$ and (b) $\mathrm{IC}_{2}$. Larger $n$ correspond to darker and thiner curves. The panels (c) and (d) correspond to the same initial conditions and show the absolute variables $|u_{1}(t)|$ for different $N=10,11,\ldots,17$.
@@
The inset of Fig. 8 shows the graphs of $\log\log\left(\|\delta u(t)\|/3\varepsilon\right)$ as functions of $N$, proposing that the growth of perturbations in the RG dynamics is double exponential in $N$. Since $l_{N}=2^{-N}$, such a double exponential function of $N$ would be exponential in a negative power of the regularization scale $l_{N}$. This relation connects our observations to earlier phenomenological results in the turbulence theory relating positive Lyapunov exponents with a dissipative scale; see e.g. [38].
@@ 8 Discussion
We study initial boundary value problems (IBVP) for ideal and regularized shell models. In general, each regularized IBVP can be represented by a flow map that contains solutions for all initial and boundary conditions. We show how the scale invariance of an ideal system defines a renormalization group (RG) operator in the space of flow maps. This operator acts by biasing the regularization towards smaller scales and introducing ideal interactions at the freed largest scale. In this way, we establish the correspondence between the vanishing regularization limit and the RG dynamics in the space of flow maps. Thereby, the limiting flow map solving the ideal IBVP is associated with the RG attractor, and the convergence is controlled asymptotically by the leading eigenmode of the linearized RG operator. Since the RG operator is defined by the ideal system (with all information on a specific regularization contained in flow maps), this yields a qualitative explanation of the universality in the regularization process. This method is first applied to the dyadic shell model, where the RG dynamics has a fixed-point attractor. Using the Gledzer-type shell model, we demonstrated a different situation, when the RG attractor is quasiperiodic. This example reveals a more sophisticated but still universal behavior in the limit of vanishing regularization.
@@
The developed RG approach is potentially applicable to realistic models, such as the Burgers and Navier-Stokes equations. Although obtaining any rigorous results using the RG approach would be very difficult, its predictions are non-trivial and can be verified numerically. This refers both to the universal limiting solutions and universal form of deviations, as we demonstrated with the dyadic model. We emphasize that our RG approach is qualitative and aims to explain the universal features of regularizations rather than to calculate specific numerical values. Another promising direction is the (potentially rigorous) application of our RG formalism to even simpler toy models such as fractal lattice models [35, 34] and point singularities in low-dimensional ordinary differential equations [12, 13, 16]. It would also be interesting to understand the relation of our RG formalism to the hidden scale invariance formulated both for shell models [31] and Navier–Stokes system [32, 36]. In particular, the extended form of the hidden symmetry [33] is limited to canonical regularizations for similar reasons, predicting the hidden self-similar statistics at small scales in both inertial and dissipative ranges. This analysis may be useful for extending the RG approach to non-canonical (but physically motivated) regularizations, following the auxiliary constructions of Sections 5 and 6.3.
In the final part of the paper we tested the Sabra shell model. This model is known to be spontaneously stochastic in the formulation that includes both viscosity and small-scale noise [29, 1]. The spontaneous stochasticity means that solutions remain stochastic in the limit when both viscous and noise terms vanish; similar behavior was reported in other shell models as well [28, 30, 3]. In this paper, we demonstrated that the RG dynamics of the Sabra model is chaotic. These two properties, the spontaneous stochasticity and the chaotic RG attractor, are related as we demonstrated in [34] for discrete space-time models. For the Sabra model, this relation will be studied in the forthcoming paper.
@@ 9.3 RG eigenvalue $\rho=-1/2$
Acknowledgments. This work was supported by CNPq grant 308721/2021-7 and FAPERJ grant E-26/201.054/2022.

### v2

@@ 6.1 RG formalism
The RG formalism is defined by the symmetries and couplings of the ideal IBVP. Similarly to the dyadic model, the symmetries are the time and space scalings. The time scaling symmetry has the same form (2.7). For the boundary condition (6.2), the space scaling symmetry is formulated as
$\tilde{u}(t)=S^{+}u(t),\quad\tilde{a}=S^{+}a,\quad\tilde{b}(t)=(\lambda b_{0}(t),\,\lambda u_{1}(t)).$ (6.3)
The RG formalism is introduced in the same way as in Section 3; see Definitions 1 and 2. Here Eqs. (3.3) and (3.4) are adapted to coupling function (6.1) and boundary condition (6.2) as
$\frac{du_{1}}{dt}=k_{1}f(b_{1},b_{0},u_{1},u_{2},u_{3}),\quad u_{1}(0)=a_{1},$ (6.4)
where $u_{2}(t)=\tilde{u}_{1}(t)/\lambda$ and $u_{3}(t)=\tilde{u}_{2}(t)/\lambda$ with
$\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b}),\quad\tilde{a}=S^{+}a;\quad\tilde{b}(t)=\big{(}\lambda b_{0}(t),\,\lambda u_{1}(t)\big{)}.$ (6.5)
[…]
@@ 6.2 RG attractor
[…]
@@
Let us consider regularizations (3.8) with $J=3$ and modified dissipative terms $c_{n}|u_{n}|u_{n}$, which are multiplied by positive coefficients $c_{n}$. We perform 100100100 simulations for each $N=40,\ldots,80$ by choosing $c_{n}$ randomly from the interval [0,3][0,3][ 0 , 3 ]. The resulting values $(u_{1},u_{2})$ are presented by black dots in Fig. 5, which form closed curves in the phase space. According to Eq. (3.9), each random choice of regularization corresponds to a specific RG sequence $\mathcal{R}^{N}[\Phi]$ starting from a different initial flow map $\Phi$. Hence, the black dots forming a curve in Fig. 5 determine the RG attractor probed by 100 different initial conditions. Similar curves appear for any pair of shell variables. The one-dimensional asymptotic structure is also apparent in Figs. 4(c,d), where the data from Fig. 5 is added at final times.
@@
Our numerical results suggest that the attractor of the RG dynamics is a closed curve $\Phi^{(\theta)}$ in the space of flow maps parametrized by points of a circle, $\theta\in S^{1}$. In the theory of dynamical systems, such attractors appear, e.g., in the Hopf bifurcation of maps [23]. Note also that the folds and self-intersections of the invariant curves in Fig. 5 are a consequence of the two-dimensional projection; examination of the three-dimensional graphs in the space $(u_{1},u_{2},u_{3})$ does not reveal any singularities. This attractor leads to a more sophisticated (than in the fixed-point case) but still universal description of the vanishing regularization limit. Namely, within the basin of attraction, $\Phi\in\mathcal{B}(\Phi^{(\theta)})$, the RG dynamics takes the asymptotic form
@@
where $\Phi^{(\theta_{N})}$ are flow maps of the attracting invariant curve. The phases $\theta_{N}$ depend on the initial flow map $\Phi$, and are related by the circle map $\mathcal{R}_{\theta}:\theta_{N}\mapsto\theta_{N+1}$ induced by the RG dynamics on the attractor. In other words, for any canonical regularization from the basin of attraction, regularized solutions $u(t)=\mathcal{R}^{N}[\Phi]_{t}(a,b)$ approach asymptotically a universal (i.e., independent of regularization) family of solutions $\Phi^{(\theta)}_{t}(a,b)$. The latter is a one-parameter family of solutions of the ideal IBVP.
@@ 6.3 Inviscid limit
Here the parameter $\theta$, which selects a map from the RG attractor, depends on $(N,\beta)$, which, in turn, are functions of viscosity, time, initial and boundary conditions. This dependence implies that the universal infinitesimal-time relation (6.7) does not extend to finite times, unlike in the case of canonical regularizations (6.6).
Figure 6: (a) Evolutions of the first two variables $(u_{1},u_{2})$ for the viscous model with the second initial condition $\mathrm{IC}_{2}$. Different curves correspond to different viscosities logarithmically spaced in the interval $1.65\times 10^{-10}\leq\nu\leq 10^{-6}$. (b) Terminal points (at time $t=0.5$) of the viscous solutions from the left panel are compared with the data from Fig. 5(b) featuring the closed-curve RG attractor.
@@ 7 Chaotic RG dynamics in the Sabra model
The RG formalism for the Sabra model is developed in the same way as in Sections 3 and 6. We learned in Section 3 that the RG operator is defined unambiguously if it is restricted to canonical flow maps, which are maximally symmetric. The latter means that the flow maps are invariant with respect to all symmetries except the space scaling (6.3). As observed in the example of time-scale invariance broken by viscous regularization, this introduces ambiguity in the definition of the RG operator (see Section 3) and also affects the universality of the limiting behavior (see Section 5). We expect that these features may extend to any additional symmetry of the system. In the Sabra model, such additional symmetries are the time scaling and phase symmetries. From now on we assume that all regularizations under consideration are invariant with respect to the phase symmetry (7.2). For a flow map $\Phi$ this condition means that the regularized solutions $u(t)=\Phi_{t}(a,b)$ and $\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b})$ are related by Eq. (7.2). One can see that the viscous model (2.8) is phase invariant, but not invariant to the time scaling.
[…]
@@
The inset of Fig. 8 shows the graphs of $\log\log\left(\|\delta u(t)\|/3\varepsilon\right)$ as functions of $N$, proposing that the growth of perturbations in the RG dynamics is double exponential in $N$. Since $l_{N}=2^{-N}$, such a double exponential function of $N$ would be exponential in a negative power of the regularization scale $l_{N}$. This relation connects our observations to earlier phenomenological results in the turbulence theory relating positive Lyapunov exponents with a dissipative scale; see e.g. [41]. A similar double-exponential growth of perturbations in the RG dynamics was also observed in discrete-time models [38, 36].
@@ 8 Discussion
We study initial boundary value problems (IBVP) for ideal and regularized shell models. In general, each regularized IBVP can be represented by a flow map that contains solutions for all initial and boundary conditions. We show how the scale invariance of an ideal system defines a renormalization group (RG) operator in the space of flow maps. This operator acts by biasing the regularization towards smaller scales and introducing ideal interactions at the freed largest scale. In this way, we establish the correspondence between the vanishing regularization limit and the RG dynamics in the space of flow maps. Thereby, the limiting flow map solving the ideal IBVP is associated with the RG attractor, and the convergence is controlled asymptotically by the leading eigenmode of the linearized RG operator. Since the RG operator is defined by the ideal system (with all information on a specific regularization contained in flow maps), this yields a qualitative explanation of the universality in the regularization process. This method is first applied to the dyadic shell model, where the RG dynamics has a fixed-point attractor. Using the Gledzer-type shell model, we demonstrated a different situation, when the RG attractor is represented by an invariant closed curve. This example reveals a more sophisticated but still universal behavior in the limit of vanishing regularization. The RG attractor represented by an invariant closed curve may also be related to the phenomenon of spontaneous stochasticity when small-scale noise is added to the equations of motion, as studied previously in a similar Gledzer-type model [30].
@@
The developed RG approach is potentially applicable to realistic models, such as the Burgers and Navier-Stokes equations. Although obtaining any rigorous results using the RG approach would be very difficult, its predictions are non-trivial and can be verified numerically. This refers both to the universal limiting solutions and universal form of deviations, as we demonstrated with the dyadic model. We emphasize that our RG approach is qualitative and aims to explain the universal features of regularizations rather than to calculate specific numerical values. Another promising direction is the (potentially rigorous) application of our RG formalism to even simpler toy models such as fractal lattice models [38, 37, 36] and point singularities in low-dimensional ordinary differential equations [12, 13, 16]. It would also be interesting to understand the relation of our RG formalism to the hidden scale invariance formulated both for shell models [33] and Navier–Stokes system [34, 39]. In particular, the extended form of the hidden symmetry [35, 27] is limited to canonical regularizations for similar reasons, predicting the hidden self-similar statistics at small scales in both inertial and dissipative ranges. This analysis may be useful for extending the RG approach to non-canonical (but physically motivated) regularizations, following the auxiliary constructions of Sections 5 and 6.3.
In the final part of the paper we tested the Sabra shell model. This model is known to be spontaneously stochastic in the formulation that includes both viscosity and small-scale noise [31, 1]. The spontaneous stochasticity means that solutions remain stochastic in the limit when both viscous and noise terms vanish; similar behavior was reported in other shell models as well [30, 32, 3]. In this paper, we demonstrated that the RG dynamics of the Sabra model is chaotic. These two properties, the spontaneous stochasticity and the chaotic RG attractor, are related as we demonstrated in [37, 36] for discrete space-time models. For the Sabra model, this relation will be studied in the forthcoming paper.
@@ 9.3 RG eigenvalue $\rho=-1/2$
Acknowledgments. This work was supported by CNPq grant 308721/2021-7, FAPERJ grant E-26/201.054/2022 and CAPES MATH-AmSud project CHA2MAN.
Data Accessibility. Data available on reasonable request.